OpenStudy (stefrheart):

Will give medal. The line (3-2k)x=7y=13 has a slope of -5/3. Find K.

OpenStudy (triciaal):

verify your post it has a typo = vs +

OpenStudy (stefrheart):

no it is an equal sign

OpenStudy (triciaal):

i got it

OpenStudy (triciaal):

standard form the first one is the +

OpenStudy (triciaal):

rewrite the equation to put in slope-intercept form given ax + by + c = 0 rewrite to y = m x + b the slope m = -5/3 can you finish from here?

OpenStudy (stefrheart):

i have a question how can you have the slope if it becomes 7y=-3(3-2k)x+13

OpenStudy (triciaal):

the value of the slope was given. are you saying the original has 2 =?

OpenStudy (stefrheart):

no it did not i looked at the wrong equals sign sorry my mistake

OpenStudy (triciaal):

so would you agree that if we have the line in slope-intercept form we can equate the expression for the slope with the given value for the slope and find the value of k?

OpenStudy (stefrheart):

yes i agree

OpenStudy (triciaal):

do you know how to use the distributive property and to manipulate fractions? equate the slope expression to -5/3 and find k I have k = -4/9

OpenStudy (stefrheart):

let me see and i know distributive not manipulating fractions

OpenStudy (triciaal):

you can cross-multiply

OpenStudy (triciaal):